When a 0.5 kg mass is hang on a spring. It stretches 0.25 mm.

A. What is the spring constant?
B. How much would a 2 kg mass stretch the string?
1
Expert's answer
2018-11-22T13:18:10-0500

a) We can find the spring constant from the Hooke’s law:

F=kx,

here, F=mg is the force acting on the spring, m is the mass of the spring, g=9.8 m⁄s^2 is the acceleration due to gravity, k is the spring constant and x is the elongation of the spring.

Then, from this equation we can find the spring constant:

mg=kx,

k=mg/x=(0.5 kg∙9.8 m/s^2 )/(2.5∙〖10〗^(-4) m)=19600 N/m.

b) As we know the spring constant, we can find the elongation of the spring from the same formula:

mg=kx,

x=mg/k=(2 kg∙9.8 m/s^2 )/(19600 N/m)=0.001 m=1.0 mm.